# A Distributed Privacy Preserving Model for the Detection of Alzheimer’s Disease

### v3

@@ 5 Discussion
Firstly, the VFL model could be trained for the diagnosis of AD, with an accuracy rate of 82.9%. This diagnostic accuracy rate is well within the results of other investigators who applied HFL for AD detection and reported accuracy rates between 62% and 89% 18; 21; 22. Despite the accuracy rate obtained, one limitation of the study is OASIS-2 dataset’s small sample size. The OASIS-2 dataset only had 150 patients, which made it challenging to train a robust neural network. Splitting the train, validation, and test sets on the imaging session or MRI image instead of on the patient axis may have yielded a higher accuracy rate. However, such a strategy also could have resulted in an over-fit because of the small sample size. Moving forward, the OASIS-3 dataset 41 consists of 1,379 patients. Additionally, more data has been included– such as both amyloid and metabolic positron emission tomography scans, neuropsychiatric testing, genetic testing, and other clinical data. Future studies, therefore, could use OASIS-3 for multimodal training of a VFL architecture for AD detection. OASIS-3 training of a VFL archtitecture most likely would produce a much more robust model trained on more modalities of data. Although training on a wide array of different datasets– as found in OASIS-3– should increase diagnostic accuracy, the availability of more data in a clinical setting could bring about further challenges if used to train the herein described VFL model. For example, the assumption that all of the demographic features rely on one federate and the image data rely on another might neither be realistic or feasible. To obviate this hurdle, one option would be to apply the Cascade Vertical Federated Learning (CVFL) framework 42 . CVFL decomposes the neural network into subnetworks, thereby allowing the utilization of partitioned labels for privacy-preserving training. CVFL allows for VFL training when some of the labels are distributed horizontally. Such a framework is advantageous particularly in scenarios where features and labels are distributed among different parties, Therefore, CVFL aligns with most clinical settings, in which data and labels often are distributed across multiple federates. The potential implementation of CVFL in future iterations of our VFL research promises a more robust and clinically relevant approach to federated learning within the healthcare domain.
@@
Secondly, by using fully homomorphic encryption to prevent the leakage of model weights, my VFL model is privacy-preserving. Since federated learning does not require sending personal health information to the central server, it inherently is HIPAA compliant. The importance of security and privacy protections is a costly concern. With regard to preserving personal health information, in 2019, it was reported that the total cost of healthcare sector data breaches totaled \$8 million dollars, or \$400 dollars per each medical record. In the first 6 months of 2024, there already have been 340 data breaches, affecting 42,250,702 individuals 43; 44. By the end of this year, a total cost of 11 million dollars or 500 dollars per medical record could be anticipated. Thus, DML must have safeguard around PHI, and– as discussed above– all federated learning frameworks inherently would protect PHI.

### v4

@@ 5 Discussion
Firstly, the VFL model could be trained for the diagnosis of AD, with an accuracy rate of 82.9%. This diagnostic accuracy rate is well within the results of other investigators who applied HFL for AD detection and reported accuracy rates between 62% and 89% [18, 21, 22]. Despite the accuracy rate obtained, one limitation of the study is OASIS-2 dataset’s small sample size. The OASIS-2 dataset only had 150 patients, which made it challenging to train a robust neural network. Splitting the train, validation, and test sets on the imaging session or MRI image instead of on the patient axis may have yielded a higher accuracy rate. However, such a strategy also could have resulted in an over-fit because of the small sample size. Moving forward, the OASIS-3 dataset [41] consists of 1,379 patients. Additionally, more data has been included– such as both amyloid and metabolic positron emission tomography scans, neuropsychiatric testing, genetic testing, and other clinical data. Future studies, therefore, could use OASIS-3 for multimodal training of a VFL architecture for AD detection. OASIS-3 training of a VFL archtitecture most likely would produce a much more robust model trained on more modalities of data. Although training on a wide array of different datasets– as found in OASIS-3– should increase diagnostic accuracy, the availability of more data in a clinical setting could bring about further challenges if used to train the herein described VFL model. For example, the assumption that all of the demographic features rely on one federate and the image data rely on another might neither be realistic or feasible. To obviate this hurdle, one option would be to apply the Cascade Vertical Federated Learning (CVFL) framework [42] . CVFL decomposes the neural network into subnetworks, thereby allowing the utilization of partitioned labels for privacy-preserving training. CVFL allows for VFL training when some of the labels are distributed horizontally. Such a framework is advantageous particularly in scenarios where features and labels are distributed among different parties, Therefore, CVFL aligns with most clinical settings, in which data and labels often are distributed across multiple federates. The potential implementation of CVFL in future iterations of our VFL research promises a more robust and clinically relevant approach to federated learning within the healthcare domain.
@@
Secondly, by using fully homomorphic encryption to prevent the leakage of model weights, my VFL model is privacy-preserving. Since federated learning does not require sending personal health information to the central server, it inherently is HIPAA compliant. The importance of security and privacy protections is a costly concern. With regard to preserving personal health information, in 2019, it was reported that the total cost of healthcare sector data breaches totaled \$8 million dollars, or \$400 dollars per each medical record. In the first 6 months of 2024, there already have been 340 data breaches, affecting 42,250,702 individuals [43, 44]. By the end of this year, a total cost of 11 million dollars or 500 dollars per medical record could be anticipated. Thus, DML must have safeguard around PHI, and– as discussed above– all federated learning frameworks inherently would protect PHI.